OpenStudy (anonymous):

Write the following as an exponential expression. 5root of x^3

OpenStudy (anonymous):

\[\sqrt[5]{x ^{3}}\]

OpenStudy (anonymous):

is it \[x^{3/5}\]

OpenStudy (angela210793):

=x^(3/5)

OpenStudy (anonymous):

\[\sqrt[5]{x^2}\] \[(x^2)^\frac{1}{5}\]

OpenStudy (anonymous):

\[(x^3)^\frac{1}{5}\] \[(x)^\frac{3}{5}\]